OpenStudy (caitlinnr14):

3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (caitlinnr14):

im confused

jimthompson5910 (jim_thompson5910):

so you're completely stuck and haven't gotten anywhere?

OpenStudy (caitlinnr14):

nope

jimthompson5910 (jim_thompson5910):

the Symmetric Property is the idea that if x = y, then y = x so that answer choice is eliminated since this isn't what we want

jimthompson5910 (jim_thompson5910):

what is the Commutative Property of Addition ?

OpenStudy (caitlinnr14):

addends can be added in any order and the sum is always the same.

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

basically the idea that x+y = y+x for any real numbers x and y

jimthompson5910 (jim_thompson5910):

so that's out too

jimthompson5910 (jim_thompson5910):

how about the Associative Property of Addition what does that say?

OpenStudy (caitlinnr14):

it doesn't matter how you add them, the totals are always same.

jimthompson5910 (jim_thompson5910):

close, more like how the grouping doesn't affect the answer

OpenStudy (caitlinnr14):

ok

jimthompson5910 (jim_thompson5910):

so a+(b+c) = (a+b)+c for any real numbers a, b, c

jimthompson5910 (jim_thompson5910):

so that choice is out

jimthompson5910 (jim_thompson5910):

the only thing left is the distributive property

jimthompson5910 (jim_thompson5910):

and it turns out that the distributive property is x(y+z) = x*y + x*z which is exactly what is used in the original problem

OpenStudy (caitlinnr14):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome